# MOTGNN: Interpretable Graph Neural Networks for Multi-Omics Disease Classification

### v1

@@ 5. Results and Discussion
Figure 5 visually compares model performance across all datasets and evaluation metrics, highlighting MOTGNN’s consistent advantage. Although some confidence intervals slightly overlap, MOTGNN consistently achieves higher mean scores. The most significant gaps appear in F1-scores for COADREAD, where MOTGNN (87.2%) far exceeds RF (33.4%), with non-overlapping confidence intervals. The figure also illustrates that, despite XGBoost’s strong performance in COADREAD (91% accuracy), its accuracy drops considerably in LGG and STAD, underscoring its lack of robustness across datasets. Figure 6 highlights model robustness on the imbalanced COADREAD dataset (class ratio 254:78) by comparing F1-scores over 20 independent experiments. MOTGNN achieves the highest median F1-score (0.871), outperforming XGBoost (0.794), DFN (0.771), GCN (0.735), and RF (0.349). The violin plot reveals a tight distribution around MOTGNN’s median, indicating low variability. In contrast, XGBoost, DFN, and GCN show wider interquartile ranges, while RF exhibits low performance and high variance. These results underscore MOTGNN’s strength in handling class imbalance, a key challenge in biomedical data analysis.
@@
Table 6 presents the top 30 most important biomarkers identified by MOTGNN for the COADREAD dataset (Colorectal Adenocarcinoma and Rectal Adenocarcinoma), illustrating the feature-level interpretability of our framework. COADREAD encompasses colorectal cancer in a broader context and is commonly analyzed jointly, as colon and rectal adenocarcinomas share overlapping molecular signatures and clinical features. This dataset was selected due to its well-characterized molecular landscape and the strong classification performance achieved by our model. The same feature-ranking procedure can also be applied to the other two cancer types: STAD (Stomach Adenocarcinoma) and LGG (Lower Grade Glioma). Among the top-ranked features, Secreted Frizzled-Related Protein 4 (SFRP4) is a well-documented biomarker associated with colorectal cancer (Huang et al., 2010; Yu et al., 2022; Frank et al., 2010; Liu et al., 2020). SFRP4 is overexpressed in tumor tissues, and its promoter hypermethylation has been observed as an early event in cancer progression. Functionally, it acts as a modulator of Wnt signaling, thereby influencing cell proliferation and tumor growth. While our current study focuses on computational identification through feature importance scores, biological validation remains an important next step. These selected biomarkers, including SFRP4, serve as potential candidates for further investigation and may contribute to discovering novel diagnostic or therapeutic strategies for diseases.
Table 6. Top 30 biomarkers identified by MOTGNN for COADREAD by omics type (methylation, mRNA, and miRNA)
@@
Note: The COADREAD dataset refers to Colorectal Adenocarcinoma and Rectal Adenocarcinoma (See data details in Table 2). Biomarkers are ranked and selected from three omics types: DNA methylation (meth), gene expression (mRNA), and microRNA expression (miRNA).
Table 7 reveals the computational trade-offs associated with these performance improvements. MOTGNN requires the longest training times, averaging 1.75 minutes for COADREAD and up to 2.39 minutes for LGG, compared to GCN’s 0.59-0.94 minutes, DFN’s 0.3-0.39 minutes, XGBoost’s 0.16-0.25 minutes, and RF’s remarkably fast 0.05 minutes per run. However, these computational costs remain practical for real-world applications, especially considering the analysis was conducted using only four CPUs per job. The increased runtime is justified by MOTGNN’s superior and stable performance across diverse datasets and metrics.
Table 7. Computation time comparison across models and datasets (in minutes).
@@ 6. Conclusion
We present MOTGNN, a novel graph neural network framework for multi-omics data integration and disease classification. By leveraging XGBoost for omics-specific supervised graph construction and modality-specific graph neural networks for latent representation learning, MOTGNN effectively captures complex relationships among features that are often overlooked by traditional methods. Applied to real-world cancer datasets, MOTGNN consistently outperforms baseline models across evaluation metrics, demonstrating its strength in integrating heterogeneous omics data and improving predictive performance. In particular, MOTGNN maintains strong robustness under class imbalance, a common challenge in biomedical data analysis. In addition to its predictive power, MOTGNN provides interpretability at both the feature and omics levels. Feature-level importance scores enable the identification of top-ranked biomarkers, which may serve as promising candidates for further biological investigation. Meanwhile, omics-level contributions reveal the relative informativeness of each data modality (e.g., methylation, mRNA, or miRNA), offering valuable insight into their role across different disease types. Future extensions of the MOTGNN framework may include incorporating additional omics modalities, additional cancer types and non-cancer diseases, or adapting the architecture for more complex prediction tasks, such as multi-class classification or survival analysis. Further enhancing interpretability will also be a key direction, especially to better support clinical and biological decision-making.
@@
The harmonized datasets used in our study and the code for our model are available upon request or publication.
@@ Acknowledgements.
This work is partially supported by the National Institute of General Medical Sciences of the National Institutes of Health (NIH/NIGMS) under Award P20GM104420. Any findings, comments, and conclusions expressed in this paper do not reflect the views of NIH.

### v2

@@ 4. Results and Discussion
Figure 4 compares model performance across all datasets and evaluation metrics, highlighting MOTGNN’s consistent advantage over both baseline and ablation models. Although some confidence intervals slightly overlap, MOTGNN consistently attains the highest mean scores across accuracy, ROC-AUC, and F1-score. The largest performance gaps appear in F1-scores for COADREAD, where MOTGNN (87.2%) far exceeds RF (33.4%) and outperforms all other methods, including $\mathrm{MOTGNN}{\text{gcn}}$ and $\mathrm{MOTGNN}{\text{rf}}$. The figure also shows that, despite XGBoost’s strong performance in COADREAD (91% accuracy), its accuracy drops markedly in LGG and STAD, underscoring limited robustness across datasets. Figure 5 further demonstrates model robustness on the imbalanced COADREAD dataset (class ratio 254:78) by comparing F1-scores over 20 independent runs. MOTGNN achieves the highest median F1-score (0.871), outperforming XGBoost (0.794), DFN (0.771), GCN (0.735), RF (0.349), and both ablation variants $\mathrm{MOTGNN}{\text{gcn}}$ (0.862) and $\mathrm{MOTGNN}{\text{rf}}$ (0.800). The violin plots reveal a compact distribution around MOTGNN’s median, indicating low variability and stable performance. In contrast, XGBoost, DFN, and GCN display wider interquartile ranges, while RF shows low performance and high variance. The ablation variants outperform their respective baselines but remain below the full MOTGNN, reinforcing that both the GEDFN backbone and XGBoost-based graph generation contribute to the model’s overall strength. These results underscore the robustness of MOTGNN under class imbalance, a key challenge in biomedical data analysis.
Graph Importance and Omics-Level Interpretability. Figure 6 displays the relative graph importance for the three constructed graphs. The observed edge-to-node ratios (ranging from approximately 2.1 to 2.8, as shown in Table 4) indicate that all graphs are sparse. Among them, the miRNA graph ($G_{3}$) shows the highest edge density (2.35-2.79) but the lowest relative importance (24-29%), suggesting that it contributes less to classification performance despite being more connected. In COADREAD, the mRNA graph ($G_{2}$) contributes most strongly, with an importance of roughly 42%. In LGG and STAD, methylation ($G_{1}$) and mRNA ($G_{2}$) contribute comparably, ranging from 35% to 37%. The consistently high methylation importance across datasets (approximately 34-37%) underscores its robustness as a discriminative biomarker, whereas the variability of mRNA importance reflects tissue-specific gene expression dynamics. These results demonstrate that MOTGNN effectively identifies the most informative omics modalities while maintaining computational efficiency through sparse graph construction. The resulting edge counts remain tractable, supporting scalability to larger or clinically oriented datasets.
@@
Biomarker Selection and Feature-Level Interpretability. Table 6 lists the top 30 most important biomarkers identified by MOTGNN for the COADREAD dataset, illustrating the feature-level interpretability of our framework. COADREAD encompasses both colorectal and rectal adenocarcinomas, which are often analyzed jointly because of their overlapping molecular signatures and clinical profiles. This dataset was selected for interpretability analysis due to its well-characterized molecular landscape and the strong classification performance achieved by MOTGNN. The same feature-ranking procedure can also be applied to the other two cancer types: STAD (Stomach Adenocarcinoma) and LGG (Lower Grade Glioma). Among the top-ranked features identified for COADREAD, Secreted Frizzled-Related Protein 4 (SFRP4) stands out as a well-documented biomarker associated with colorectal cancer (Huang et al., 2010; Frank et al., 2010; Liu et al., 2020; Yu et al., 2022). Multiple studies have reported that SFRP4 is overexpressed in tumor tissues, and its promoter hypermethylation occurs as an early event in cancer progression. Functionally, SFRP4 acts as a modulator of Wnt signaling, thereby influencing cell proliferation and tumor growth. While the present work focuses on computational discovery through feature-importance analysis, these findings highlight MOTGNN’s potential for guiding downstream biological validation. The selected biomarkers serve as promising candidates for further investigation and may ultimately contribute to discovering novel diagnostic or therapeutic strategies for complex diseases.
Table 6. Top 30 biomarkers identified by MOTGNN for the COADREAD dataset (Colorectal and Rectal Adenocarcinoma), grouped by omics type: DNA methylation (meth), gene expression (mRNA), and microRNA expression (miRNA). The listed biomarkers represent the biomarkers with the highest feature-importance scores derived from MOTGNN.
@@
Table 7 summarizes the computational trade-offs associated with these performance improvements. Among all models, MOTGNN requires the longest training time, averaging 1.75 minutes for COADREAD and up to 2.39 minutes for LGG, compared to GCN’s 0.59-0.94 minutes, DFN’s 0.3-0.39 minutes, XGBoost’s 0.16-0.25 minutes, and RF’s remarkably fast 0.05 minutes per run. Despite this additional cost, the runtime remains practical for real-world applications, especially considering the analysis was conducted using only four CPUs per job. The longer runtime primarily reflects the additional computational steps required for supervised graph construction and hierarchical representation learning. These added computations are justified by MOTGNN’s superior and stable performance across diverse datasets and evaluation metrics, demonstrating a favorable balance between efficiency and predictive accuracy.
Table 7. Computation time comparison of MOTGNN and baseline models (XGBoost, RF, DFN, and GCN) across COADREAD, LGG, and STAD datasets. Each experiment was conducted using only four CPUs without GPU acceleration, and the reported times are averaged over 20 independent runs.
@@ 5. Conclusion
We present MOTGNN, a novel graph neural network framework for multi-omics data integration and disease classification. By leveraging XGBoost for omics-specific supervised graph construction and modality-specific graph neural networks for latent representation learning, MOTGNN effectively captures complex relationships among features that are often overlooked by traditional methods. Applied to real-world cancer datasets, MOTGNN consistently outperforms baseline models across evaluation metrics, demonstrating its strength in integrating heterogeneous omics data and improving predictive performance. In particular, MOTGNN maintains strong robustness under class imbalance, a common challenge in biomedical data analysis. In addition to its predictive power, MOTGNN provides interpretability at both the feature and omics levels. Feature-level importance scores enable the identification of top-ranked biomarkers, which may serve as promising candidates for further biological investigation. Meanwhile, omics-level contributions reveal the relative informativeness of each data modality (e.g., methylation, mRNA, or miRNA), offering valuable insight into their roles across different disease types. Together, these interpretability results show that MOTGNN not only enhances prediction performance but also yields biologically meaningful insights that can guide downstream validation and hypothesis generation.
Future extensions of the MOTGNN framework may include incorporating additional omics modalities, expanding to other cancer types and non-cancer diseases, or adapting the architecture for more complex prediction tasks, such as multi-class classification or survival analysis. Further enhancing interpretability and computational scalability will also be important directions, particularly to better support clinical and biological decision-making. In the long term, MOTGNN’s modular design and interpretable learning process provide a flexible foundation for advancing integrative multi-omics research and enabling precision medicine applications.
@@
The original datasets are from TCGA Broad GDAC Firehose (https://gdac.broadinstitute.org/). The harmonized datasets used in our study and the code for our model are available upon publication.
@@ Acknowledgements.
This work is partially supported by the National Institute of General Medical Sciences of the National Institutes of Health (NIH/NIGMS) under Award P20GM104420.